A steel bar has a width of 10cm at 50°C. At what temperature will it fit exactly into a hole of constant width 10.005cm if coefficient of linear expansion of steel is 11*10^-6C-¹
1
Expert's answer
2019-04-08T09:36:32-0400

By the definition of the linear thermal expansion we have:


"\dfrac{\Delta L}{L} = \alpha \Delta T,""\dfrac{L - L_0}{L} = \alpha (T - T_0),"

here, "\Delta L = L - L_0" is the difference in the width of the steel bar after the change in the temperature, "L_0 = 0.1 m" is the initial width of the steel bar at temperature "T_0 = 50^{\circ}C", "L = 0.10005 m" is the width of the steel bar after the change in the temperature, "\alpha" is the coefficient of linear expansion for the steel bar, "\Delta T" is the change in temperature.

Then, from this equation we can find the temperature at what the bar will fit exactly into a hole of constant width:


"T = \dfrac{L - L_0}{\alpha L} + T_0 = \dfrac{0.10005 m - 0.1 m}{11 \cdot 10^{-6} \dfrac{1}{^{\circ}C} \cdot 0.10005 m} + 50^{\circ}C = 95.5^{\circ}C."

Answer:

"T = 95.5^{\circ}C."